[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jul 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 28 20:24:01 UTC 2017


HSU Shift Log for Fri Jul 28 2017 16:23:55 GMT-0400 (EDT)

Shift: H

Forecaster: D Zelinsky

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: Major NETAPP migration between 10:30a and 2:30p.  
All NAWIPS and ATCF workstations were restarted to ensure they are pointing  
at the new NETAPP. All operational servers also migrated and rebooted as  
necessary. Lost satellite imagery in NAWIPS for a couple hours and DAZ had  
to run the 18Z outlooks by hand on the compute server, but overall impact  
on operations was minimal. Matt S. is on call for the weekend if any  
lingering issues related to the transition come up.

Miscellaneous issues: GOES-16 meso area over Hilary continues through 0300  
UTC.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170728/fd179ede/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list